OpenStudy (anonymous):

HELPPPPPPPP!!!!!!!!!!! Estimate 3 3/4 x 5 11/12 by rounding each factor to the nearest whole number

OpenStudy (anonymous):

thank u for helping jhonny

jhonyy9 (jhonyy9):

hi just for an example 1 1/2 = (1*2+1)/2 = 3/2 hope you understand the method clearly

OpenStudy (anonymous):

hmm...

jhonyy9 (jhonyy9):

so than 3 3/4 = (3*4+3)/4 = (12 +3)/4 = 15/4 so the second try calculi it you

OpenStudy (anonymous):

aha....

jhonyy9 (jhonyy9):

so than 5 11/12 = ?

OpenStudy (anonymous):

dude i have no idea how to

OpenStudy (anonymous):

@hero! long time no see!!!!!

jhonyy9 (jhonyy9):

multiple the fisrt number what is in this case 5 by denominator what is 12 so than will be 5*12 = 60 and assum to this number 11 = 71 /12

hero (hero):

Make a number line: